OpenStudy (anonymous):

What is the simplified form of each expression? √169

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Because 13^2 = 13*13 = 169

OpenStudy (anonymous):

Simplify the expression. (8 + 7a) + 4